OpenStudy (anonymous):

I have a series of cross products that I'm trying to condense down as follows...

OpenStudy (anonymous):

\[ma = m r'\times (Curl (\Omega \times r))+ Gradient(m/2(\Omega \times r)^2)\]

OpenStudy (anonymous):

into \[ma = F + 2m r' \times \Omega + m ((\Omega \times r) \times \Omega)\]

OpenStudy (anonymous):

My vector calc is lacking and I was wondering if I could get some help besides the BAC-CAB rules. Or some intuition that might be really helpful to not do pages of algebra